6.The current one-year Treasury bill rate is 4.2%. The expected one-year rate, one year from now is 5.1%. According to the pure expectations theory, what should be today’s rate for a two-year Treasury security.

Select one:
a. 4.200%
b. 4.649%
c. 5.100%
d. 3.050%
e. 4.973%

15,Until the 1960's, limits on the interest rates that banks could pay to depositors had virtually no impact on the ability of banks to compete with other financial institutions to obtain funds because:

Select one:
a. Market interest rates stayed below the ceiling rate.
b. Market interest rates stayed above the ceiling rate.
c. Market interest rates and the ceiling rate stayed the same.
d. Market interest rates did not change.
e. Market interest rates did not change.

38.A market is price efficient if:

Select one:
a. It is riskless.
b. It offers investors reasonably priced services related to buying and selling.
c. Asset prices are as high as possible
d. There are no transaction costs and taxes.
e. Prices reflect all information that is relevant to the valuation of securities.
Answer #1

6.Answer is B

S1 = 4.2%

1f1 = 5.1%

S2 = ??

So as per expectation theory

(1+S2)^2 = (1+S1)*(1+1f1)

(1+S2)^2 = (1.042)*(1.051)

S2 = 4.649%

38. Answer is E

If the prices reflects all the relevant information, then it is price efficient

I am sorry but i am not sure about 15, so do not want to give wrong answer. I Think other 2 are correct

Request you to please do not rate thumb down based on that